OpenStudy (anonymous):

Calculate the length of the path over the given interval. x(t)=8t^(2), y(t)=6t^(2)-1. 0 less or equal t more or equal than 4

OpenStudy (loser66):

parametric equation of the path is x(t) = 8t^2; y(t) = 6t^2 -1 ; 0<= t<= 4

OpenStudy (anonymous):

It's a webwork question. I'm not sure what you mean. Yeah that is the parametric equation

OpenStudy (anonymous):

I think it's arc length.

OpenStudy (loser66):

oh, I see, it's arc length from cal2

OpenStudy (anonymous):

yeah with integration.

OpenStudy (loser66):

@zepdrix

zepdrix (zepdrix):

grr i can't remember D: Lemme think about it.

OpenStudy (loser66):

hahahaa.... relax....

zepdrix (zepdrix):

So the length of the path, arc length, will be,\[\large L=\int\limits ds=\int\limits \sqrt{1+\left(\frac{dy}{dx}\right)^2}dx\] Oh but we're dealing with parametric equations, so the formula we want to use is,\[\large L=\int\limits\limits \sqrt{\left(\frac{dx}{dt}\right)^2+\left(\frac{dx}{dt}\right)^2} dt\]

zepdrix (zepdrix):

The limits on our integral will be the interval of t,\[\large L=\int\limits_0^4 \sqrt{\left(\frac{dx}{dt}\right)^2+\left(\frac{dy}{dt}\right)^2}\]

zepdrix (zepdrix):

So this shouldn't be too bad, we just need to find a couple of derivatives, then integrate. Can you do the first one? \(\large x=8t^2\) \(\large \dfrac{dx}{dt}=?\)

OpenStudy (anonymous):

is that 16t?

zepdrix (zepdrix):

\[\large L=\int\limits\limits\limits_0^4 \sqrt{\left(16t\right)^2+\left(\frac{dy}{dt}\right)^2}dt\]Ok good.

zepdrix (zepdrix):

Understand what I did? I skipped a step there. I plugged the 16t in for our dx/dt.

OpenStudy (anonymous):

yeah

zepdrix (zepdrix):

cool, what do you get for the other derivative? c:

OpenStudy (anonymous):

12t

zepdrix (zepdrix):

\[\large L=\int\limits\limits\limits\limits_0^4 \sqrt{\left(16t\right)^2+\left(12t\right)^2}dt\] Ok good. So we've successfully setup our integral using the derivative pieces that we needed. Can you solve it from here? Just requires a little simplification, and an easy integral.

OpenStudy (anonymous):

Yeah. I will try. I will post if I have any problem. Thanks for the help.